A stock had returns of 29%, 14%, 23%, -8%, 9%, -14% over the last six years. What are the arithmetic and geometric means for this stock?
Answer #1

Arithemetic return = sum of return/count of return = (29+14+23-8+9-14)/6 = 8.83%

Geometric return = (product of (1+ rate) for each year)^(1/n)-1

=((1+0.29)*(1+0.14)*(1+0.23)*(1-0.08)*(1+0.09)*(1-0.14))^(1/6)-1 = 7.69%